

********************************************************************
UARS WEEKLY STATUS REPORT - 10 Dec 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM resumed operations on Monday, Dec. 6. Since then, the
experiment successfully gathered daily mid (1.1 nm) resolution,
MgII, occultation, solar pointing offset, and continuous monitoring
scans for time periods with UARS telemetry or tape recording.
A solar minimum spectrum derived from the SUSIM UV spectral
irradiance data is available from the SUSIM internet site. Details
may be found from:
http://wwwsolar.nrl.navy.mil/susim_uars_news.html
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 November 2004 (Orbit 72308, GMT Day 332) through
03 December 2004 (Orbit 72413, GMT Day 338).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 333/2321 72338 POWER OFF
PEM 335/2317 72368 POWER OFF
SUSIM 336/2234 72382 POWER OFF
SOLSTICE 336/2315 72383 FULL TIME TO DAY ONLY
A Forward-to-Reverse Yaw Around maneuver (number 135) was successfully
executed in Orbit 72393 beginning 337/ 15:19:59Z and completing at
337/ 15:57:09Z on 02 December 2004. All instruments were in their off
or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 57.34 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 25.92 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.16 and 20.04 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 12
January 2005 (DOY 012).
The beta angle ranged between the angles of 18.2 degrees to 03.7
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 23 December 2004 (DOY 358).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.95 degrees C. Additional
battery performance data for each battery on 26 November and 03
December 2004 are provided in Appendix D.
The clock error ranged from 10.9 to -09.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 01
PEM = 02
SUSIM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT called a meeting with FORMATS' software development
to discuss the file transfer issues from the UARS UPS to the UARS CMS
via FORMATS. Software development was informed of problems being seen
in the file. They will now look into their configuration tables and
scripts, to see if the cause can be isolated and/or fixed. The FOT
will call another meeting next week to review their progress.
MEETINGS
The FOT called a meeting with FORMATS and CMS software development on
30 November 2004 to discuss the SMA problems.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 January 2005 (DOY 012).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 03 December 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
72311 332/0432 04-328 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
72338 333/2321 04-329 HALOE POWER OFF
72368 335/2317 04-330 PEM POWER OFF
72379 336/1755 04-331 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
72382 336/2234 04-332 SUSIM POWER OFF
72383 336/2315 04-333 SOLSTICE FULL TIME TO DAY ONLY
72392 337/1434 04-334 MACS YAW MANEUVER
72393 337/1515 04-335 PEM YAW COMMANDS
72393 337/1515 04-334 MACS YAW MANEUVER
72393 337/1628 04-335 PEM YAW COMMANDS
72393 337/1628 04-334 MACS YAW MANUEVER
72399 338/0142 04-336 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 03 December 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 03 December 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
72311 332/0442 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
72379 336/1807 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
72399 338/0144 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 03 December 2004
26 November 2004 - GMT Day 331 Beta = 21.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.74/-3.99 +00.0/+00.0 07.0 0.00 0.0
2 22.4/ -50.4 +6.43/+5.14 +28.8/-14.4 26.2 1.04 19.3
3 11.2/ -67.2 +5.14/+3.87 +28.0/-14.8 26.1 1.04 19.0
03 December 2004 - GMT Day 338 Beta = 03.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 22.4/ -89.6 +5.46/+4.82 +27.2/-12.8 25.9 1.04 19.1
3 -11.2/-212.8 +4.50/+3.55 +26.8/-12.8 25.8 1.03 18.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 03 December 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-086 FOT 72338 334/0024 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-089 FOT 72289 330/1816 37679 TDRSS-E Data Dropout
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-086 FOT See List Nov 04 30832 Generic Late Acquisition